# ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction

### v1

@@ Abstract
Large language models (LLMs) are widely used as evaluators for open-ended tasks, while previous research has emphasized biases in LLM evaluations, the issue of non-transitivity in pairwise comparisons remains unresolved: non-transitive preferences for pairwise comparisons, where evaluators prefer A over B, B over C, but C over A. Our results suggest that low-quality training data may reduce the transitivity of preferences generated by the Evaluator LLM. To address this, We propose a graph-theoretic framework to analyze and mitigate this problem by modeling pairwise preferences as tournament graphs. We quantify non-transitivity and introduce directed graph structural entropy to measure the overall clarity of preferences. Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max), reflecting low overall clarity of preferences. To address this issue, we designed a filtering strategy, ELSPR, to eliminate preference data that induces non-transitivity, retaining only consistent and transitive preference data for model fine-tuning. Experiments demonstrate that models fine-tuned with filtered data reduce non-transitivity by 13.78% (from 64.28% to 50.50%), decrease structural entropy by 0.0879 (from 0.8113 to 0.7234), and align more closely with human evaluators (human agreement rate improves by 0.6% and Spearman correlation increases by 0.01).
ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction
[…]
@@ 1 Introduction
With the rapid advancement of large language model (LLM; OpenAI et al. 2024; Qwen Team 2025; Meta AI 2024a) technology, an increasing number of models have become available, making it essential to evaluate their capabilities for selecting the most suitable one. However, existing benchmarks such as MMLU Hendrycks et al. (2020) and HELM Liang et al. (2022) have been shown to be insufficient for capturing performance differences in open-ended tasks Zheng et al. (2024).
Given the lack of definitive answers in open-ended tasks, human expert evaluation is considered the gold standard. Yet, due to its high cost and limited scalability, the mainstream approach has shifted toward using LLM-as-a-Judge for efficient evaluation. Recent studies have demonstrated that powerful LLMs, such as GPT-4, can achieve high consistency with human judgments Zheng et al. (2024); Dubois et al. (2023). Among various approaches, pairwise comparison has emerged as the dominant paradigm, owing to its strong alignment with human preferences Samvelyan et al. (2024); Chen et al. (2024a); Li et al. (2023); Chiang et al. (2024); Liu et al. (2024); Liusie et al. (2024).
Despite these promising results, LLM-as-a-Judge still suffers from various biases—such as position, verbosity, conformity, and self-reinforcement bias—which compromise evaluation reliability Xu et al. (2024b); Zheng et al. (2024); Koo et al. (2024); Ye et al. (2025). A particularly underexplored yet critical issue is the non-transitivity of preferences generated by Evaluator LLMs—e.g., $A\succ B$, $B\succ C$, $C\succ A$, where $\succ$ denotes "is preferred to", i.e., $A\succ B$ means $A$ is preferred over $B$. An illustration is provided in Figure 1.
The first study to identify this issue Xu et al. (2025) analyzed inconsistencies in evaluation results when GPT-4 acted as the evaluator and performed pairwise comparisons using different baseline models. However, the study did not propose a method to mitigate this non-transitivity. Alarmingly, this issue is also inherited by specialized evaluators like JudgeLM, PandaLM, and Auto-J Zhu et al. (2025); Wang et al. (2024b); Li et al. (2024a), which are trained via knowledge distillation from foundation models. The distillation process may inadvertently propagate non-transitive judgment patterns to downstream evaluators.
We hypothesize that the presence of low-quality training data may impair the transitivity of the preferences generated by the Evaluator LLM. Many pairwise comparisons—particularly those from open-ended tasks—lack clear ground truth due to their subjective nature. Even human annotators often disagree, with inter-annotator agreement as low as 65.7% Li et al. (2023), indicating that such tasks inherently lack a universally accepted preference order. Consequently, training on ambiguous or low-quality comparisons can introduce or exacerbate non-transitive behavior in evaluators. This highlights the necessity of filtering unreliable data prior to model training.
In this work, we propose ELSPR, a simple yet effective method based on graph theory to assess and mitigate preference non-transitivity in Evaluator LLMs. We model multi-response pairwise comparisons as tournament graphs and filter out preference data that induce the non-transitivity of overall preferences. To further quantify the overall clarity of preference results, we introduce the use of two-dimensional structural entropy on directed graphs. Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max). Experimental results show that fine-tuning models on the filtered data leads to a 13.78% reduction in preference non-transitivity, a 0.0879 decrease in structural entropy, and improved alignment with human judgments—reflected in a 0.6% increase in human agreement rate and a 0.01 gain in Spearman correlation.
Specifically, our contributions are:
  • We introduce a graph-theoretic approach to systematically investigate the non-transitivity issue in preference data generated by Evaluator LLMs, revealing significant non-transitivity in the preference data produced by advanced LLMs.
  • We innovatively use the two-dimensional structural entropy of the directed graph to further evaluate the overall clarity of the preferences generated by large-scale LLMs.
  • We propose a effective filtering methodology that leverages tournament graphs to identify and remove preference data that induces non-transitivity, ensuring the retention of only consistent and transitive preferences. This enhances the reliability of Evaluator LLMs.
In addition, we released our code and datasets to facilitate further research into this issue [^1] https://github.com/yy0525/ELSPR.
@@ 2.1 LLM-as-a-Judge and Its Non-Transitive Preferences
The prevailing LLM-as-a-Judge system predominantly operates through pairwise comparisons, as exemplified by frameworks such as VicunaEval, AlpacaEval, and Arena-HardChiang et al. (2023); Li et al. (2023, 2024b). Specifically, these systems collect responses generated by different LLMs to a given set of questions, and use advanced LLMs as judges to evaluate the preference order between two responses, assessing the relative strengths and weaknesses of different LLMs. Recent studies highlight that even advanced models like GPT-4 exhibit significant preference non-transitivity in LLM-as-a-Judge systems, which severely undermines the reliability of preference outcomes.
Despite prior efforts exploring six prompt templates to alleviate preference non-transitivity, improvements have been limited Xu et al. (2025). To address this critical issue, we investigate the underlying conditions that give rise to non-transitivity in LLM-as-a-Judge systems and propose a novel method to mitigate such inconsistencies.
Figure 2: ELSPR(Evaluator LLM training data Self-purification non-transitive Preferences via tournament graph Reconstruction) framework overview. The collected preference data are modeled as a directed graph. Strongly Connected Components (SCCs) are identified to analyze non-transitive structures, and each SCC is reconstructed into a DAG to filter out the specific preference data responsible for the overall non-transitivity.
@@ 2.2 Data Selection for LLM Fine-Tuning
Prior works Chen et al. (2024b); Ge et al. (2024); Li et al. (2024c) have highlighted the importance of filtering high-quality subsets from original datasets. Fine-tuning with such subsets not only enhances model performance but also reduces computational costs. However, this strategy remains underexplored in the context of the LLM-as-a-Judge task. Existing studies have primarily focused on scaling model size or improving prompt engineering, overlooking the potential benefits of training data quality. Our goal is to incorporate data filtering techniques into the LLM-as-a-Judge setting. We find that fine-tuning the LLM on preference data that has been filtered to induce non-transitivity can significantly reduce preference non-transitivity and improve overall preference clarity—even outperforming the teacher model.
@@ 3 Methodology
In this section, we present a graph-theoretic formulation of the LLM-as-a-Judge task using tournament graphs to analyze pairwise preferences from Evaluator LLMs. We then introduce quality analysis criteria for training data, including non-transitivity detection via Strongly Connected Components (SCCs) and preference clarity analysis via graph entropy. Finally, we propose a data filtering method to mitigate non-transitive preferences. Figure 2 outlines the overall framework, with implementation details provided in the following subsections.
@@ 3.1 Background
The typical LLM evaluation systems rely on the transitivity assumption, that is, if the evaluation results are $A\succ B$ and $B\succ C$, then it is assumed that $A\succ C$ However, evaluations in the real world often reveal non-transitivity preference cycles ($A\succ B\succ C\succ A$), which renders traditional ranking methods ineffective Xu et al. (2025). In order to overcome this fundamental limitation, we propose a tournament-based evaluation framework grounded in directed graphs, which facilitates the application of graph theory methods for a deeper analysis of the degree of non-transitivity in preferences and the overall clarity of the preferences.
Modeling preferences generated by LLM as a tournament graph. For each question $q_{i}\in Q$, given the response set $A_{i}=\{a_{1},a_{2},...,a_{n}\}$ from $n$ LLMs, we construct a tournament graph $G_{i}=(V_{i},E_{i})$ through the following procedure:
vertices: set $V_{i}=\{v_{1},v_{2},...,v_{n}\}$ corresponds to responses $A_{i}=\{a_{1},a_{2},...,a_{n}\}$.
Edges: Defined by preferences generated by LLM:
$E_{i}=\bigcup_{\begin{subarray}{c}1\leq j,k\leq n\\ j\neq k\end{subarray}}\begin{cases}v_{k}\rightarrow v_{j},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`win'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`lose'}\end{aligned}\\[10.00002pt] v_{j}\rightarrow v_{k},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`lose'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`win'}\end{aligned}\\[10.00002pt] v_{j}\leftrightarrow v_{k},&\text{otherwise}\end{cases}$ (1)
@@
Considering the common position bias in Evaluator LLMs Wang et al. (2024a), we apply position swapping by comparing each response pair in both orders: $\mathcal{J}(a_{j},a_{k})$ and $\mathcal{J}(a_{k},a_{j})$. This ensures a more robust and balanced assessmentZheng et al. (2024). If the preferences differ across orders—indicating possible position bias—a bidirectional edge is added between the corresponding vertices to represent a ‘tie’.
3.2 Quality Analysis Criteria for LLM-as-a-Judge Training Data
Based on the directed graph representation of the tournament we constructed above (Section 3.1). We propose a quality analysis criteria for the training data of Evaluator LLM, which utilizes SCCs analysis to measure the degree of preference non-transitivity and employs two-dimensional structural entropy of directed graph structure entropy to evaluate the overall clarity of preferences.
3.2.1 Analysis of Non-Transitivity in Preferences Based on SCCs
The SCCs of a directed graph are maximal subgraphs in which, for any pair of vertices, there exists a directed path from one to the other and vice versa. This mutual reachability property aligns with the notion of preference cycles, which directly reflects non-transitivity in the preferences of the Evaluator LLM. A directed path from vertex $v_{i}$ to $v_{j}$ can be interpreted as a preference chain $a_{j}\succ\cdot\succ a_{i}$, while a reverse path from $v_{j}$ to $v_{i}$ corresponds to $a_{i}\succ\cdot\succ a_{j}$. The presence of both paths implies conflicting preferences and thus a violation of transitivity.
We use Tarjan algorithm to identify the SCCs of a directed graph. Tarjan algorithmTarjan (1972) is an efficient method for finding SCCs, based on depth-first search (DFS), with a time complexity of $O(N+M)$, where $N$ is the number of vertices and $M$ is the number of edges. Furthermore, since non-transitive relations occur between three or more vertices, we focus on components with a size greater than 2. Additionally, we filter out SCCs where there exists a bidirectional edge between every pair of vertices to eliminate cases where the relationship between vertices is a ‘tie’. For example, in the case of $(A=B),(B=C),(C=A)$, although $(A,B,C)$ forms an SCC, the preference relation is transitive. We denote the set of all SCCs in the directed graph $G$ as $SCCs(G)$,the following formula to obtain SCCs containing vertices with non-transitive relationships:
$\mathcal{S}_{\text{n-t}}=\left\{S\in SCCs(G)\,\middle|\,\begin{aligned} |S|>2\land\exists v_{j},v_{k}\in S,\\ (v_{j}\leftrightarrow v_{k})\notin E(S)\end{aligned}\right\}.$ (2)
To quantify the degree of non-transitivity in the preferences generated by the Evaluator LLM, we compute the ratio of non-transitive vertices to the total number of vertices across all directed graphs:
@@ 3.2.1 Analysis of Non-Transitivity in Preferences Based on SCCs
where the numerator represents the total number of vertices belonging to non-transitive SCCs across all questions $\mathbf{Q}$, and the denominator represents the total number of vertices across all graphs. A higher value of $\rho_{\text{non-trans}}$ indicates a greater prevalence of non-transitive preferences.
3.2.2 Analysis of Preference Clarity Based the Entropy of Directed Graph Structures
The Shannon entropyShannon (1948) is a fundamental metric in information theory used to quantify the uncertainty associated with a random variable. It represents the expected level of uncertainty in a system: the greater the uncertainty, the higher the entropy. It is defined as:
$H(X)=-\sum p(x)\log p(x),$ (4)
where $p(x)$ denotes the probability distribution of the event $x$.
@@ 3.2.2 Analysis of Preference Clarity Based the Entropy of Directed Graph Structures
Structural entropy Li and Pan (2016) extends Shannon entropy to directed graphs, providing a measure of system uncertainty and the complexity of relationships within the graph. These concepts have been widely applied across various domains Zou et al. (2024); Duan et al. (2024); Peng et al. (2024); Hou et al. (2025). We introduce the Two-Dimensional Structural Entropy of a Directed Graph to analyze the overall clarity of preferences in Evaluator LLM.
Two-Dimensional Structural Entropy of a Directed Graph: The core idea of two-dimensional structural entropy is to partition the graph’s vertices into distinct communities and compute the complexity of both intra-community and inter-community relationships. This framework effectively quantifies the overall structural uncertainty and the complexity of relationships within the graph.
In this work, we innovatively adopt SCCs as the fundamental communities, motivated by properties from directed graph theory. Within an SCC, cyclic and non-transitive preference relations exist, while between SCCs, the preference relations are strictly transitive. This distinction allows us to quantitatively assess the overall preference clarity by computing the two-dimensional structural entropy of the tournament graphs constructed by Evaluator LLM.
@@
It is important to note that SCCs composed of a single vertex inherently exhibit strict transitivity. The preference relations among such single-vertex components form a linear order that does not increase the complexity of the preference results. For example, if the original graph contains a preference $A\succ B$, the overall preference is clear. When new preferences $B\succ C$ and $A\succ C$ are added, the resulting linear order $A\succ B\succ C$ still remains clear and consistent — since the vertices $A$, $B$, and $C$ each form their own singleton SCC. This clarity reflects the inherent transitivity of singleton SCCs. Based on this observation, we exclude interactions among pure singleton SCCs when computing structural entropy. Instead, we retain only two types of interactions: (1) between singleton SCCs and multi-vertice SCCs, and (2) between multi-vertice SCCs. This selective inclusion ensures a more accurate and meaningful quantification of global structural complexity. Consequently, the closer the priority relations among vertices are to a linear order, the lower the resulting entropy.
The definition of the two-dimensional structural entropy of a directed graph based on SCC decomposition is as follows: Given a directed graph $G=(V,E)$, $n=|V|$ denote the number of vertices. Based on the SCC decomposition $SCCs(G)=\{SCC_{1},\dots,SCC_{L}\}$.
For a vertex $v\in V$, we use $d_{\text{in}}(v)$ and $d_{\text{out}}(v)$ to represent the in-degree and out-degree of $v$ in $G$, respectively. For any $SCC_{i}\subseteq SCCs$, define its volume as the total in-degree of its vertices:
$v(SCC_{i})=\sum_{v\in SCC_{i}}d_{\text{in}}(v).$ (5)
Accordingly, $v(G)$ denotes the total in-degree of the entire directed graph.
The two-dimensional structural entropy computation formula is as follows:
$H^{2}(G)=-\sum_{j=1}^{L}\frac{g_{j}}{v(G)}\log_{2}\frac{v(SCC_{j})}{v(G)}-\sum_{j=1}^{L}$ (6)
$\frac{v(SCC_{j})}{v(G)}\left(\sum_{v\in SCC_{j}}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\log_{2}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\right),$
where, $g_{j}$ denotes the total number of incoming edges to $V_{j}$ from external vertices. Specifically, it counts the edges between singleton and multi-vertice SCCs, as well as those among multi-vertice SCCs. It serves as a measure of the external influence on $V_{j}$.
Specifically, the normalized structural entropy for a graph $G$ is defined as:
$\tau(G)=\frac{H^{2}(G)}{\log n},$ (7)
where $H^{2}(G)$ represents the two-dimensional structural entropy of $G$, and $n$ denotes the number of vertices in the graph.
To evaluate the overall preference clarity of the Evaluator LLM across the set of questions $Q$, we compute the average normalized structural entropy, defined as follows:

### v3

@@ Abstract
Pairwise evaluation of large language models (LLMs) has become the dominant paradigm for benchmarking open-ended tasks, yet non-transitive preferences—where evaluators prefer A over B, B over C, but C over A—fundamentally undermine ranking reliability. We show that this critical issue stems largely from low-quality data that contains inherently ambiguous preference pairs. To address this challenge, we propose ELSPR, a principled graph-theoretic framework that models pairwise preferences as tournament graphs and systematically identifies problematic training data. ELSPR quantifies non-transitivity through strongly connected components (SCCs) analysis and measures overall preference clarity using a novel normalized directed graph structural entropy metric. Our filtering methodology selectively removes preference data that induce non-transitivity while preserving transitive preferences. Extensive experiments on the AlpacaEval benchmark demonstrate that models fine-tuned on ELSPR-filtered data achieve substantial improvements: a 13.8% reduction in non-transitivity, a 0.088 decrease in structural entropy, and significantly enhanced discriminative power in real-world evaluation systems. Human validation confirms that discarded data exhibit dramatically lower inter-annotator agreement (34.4% vs. 52.6%) and model-human consistency (51.2% vs. 80.6%) compared to cleaned data. These findings establish ELSPR as an effective data self-purification approach for developing more robust, consistent, and human-aligned LLM evaluation systems.
Code & Datasetshttps://github.com/yy0525/ELSPR
@@ 1 Introduction
With the rapid advancement of large language model (LLM; openai2024gpt4technicalreport; qwen25max; llama31) technology, an increasing number of models have become available, making it essential to evaluate their capabilities for selecting the most suitable one. However, existing benchmarks such as MMLU (hendrycks2020measuring) and HELM (liang2022holistic) have been shown to be insufficient for capturing performance differences in open-ended tasks (zheng2024judging_mt).
Given the lack of definitive answers in open-ended tasks, human expert evaluation is considered the gold standard. Yet, due to its high cost and limited scalability, the mainstream approach has shifted toward using LLM-as-a-Judge for efficient evaluation. Recent studies have demonstrated that powerful LLMs, such as GPT-4, can achieve high consistency with human judgments (zheng2024judging_mt; NEURIPS2023_5fc47800). Among various approaches, pairwise comparison has emerged as the dominant paradigm, due to its strong alignment with human preferences (NEURIPS2024_8147a43d; chen-etal-2024-humans; alpaca_eval; chiang2024chatbot; liu2024aligning; liusie-etal-2024-llm).
Despite these promising results, LLM-as-a-Judge still suffers from various biases—such as position, verbosity, conformity, selection, and self-reinforcement bias—which compromise evaluation reliability (xu-etal-2024-pride; zheng2024judging_mt; koo-etal-2024-benchmarking; ye2025justice; wei-etal-2024-unveiling; choi-etal-2025-mitigating). A particularly underexplored yet critical issue is the non-transitivity of preferences generated by evaluator LLMs (e.g., $A\succ B$, $B\succ C$, $C\succ A$), where $\succ$ denotes “is preferred to”; that is, $A\succ B$ means $A$ is preferred over $B$. An illustration is provided in Figure 1. Recent research (xu2025investigating) analyzed inconsistencies in the GPT-4 evaluation outcomes when conducting pairwise comparisons between different baseline models. Building on the widely adopted AlpacaEval framework, their findings indicate that preference non-transitivity significantly undermines the robustness of evaluation systems, leading to inconsistent model rankings under various baselines and thereby compromising the reliability of LLM-based evaluators. However, their analysis is limited to observing preference non-transitivity within triplets of samples and does not extend to larger sample sets. Moreover, they do not propose an effective method to substantially reduce the degree of non-transitivity exhibited by evaluator LLMs. These limitations highlight the need for more comprehensive investigations, particularly with larger sample sets, to further explore preference non-transitivity within LLM-based evaluation frameworks.
Worryingly, this issue is also inherited by specialized evaluators such as JudgeLM, PandaLM, and Auto-J (zhu2025judgelm; wang2024pandalm; li2024generative), which are trained through the distillation of knowledge from advanced models. The distillation process may inadvertently propagate non-transitive judgment patterns to downstream evaluators.
We hypothesize that the presence of low-quality training data may impair the transitivity of the preferences generated by the evaluator LLM. Many pairwise comparisons, particularly those from open-ended tasks, lack definitive ground truth due to their inherent subjectivity. Human annotators frequently demonstrate significant disagreement, with empirical studies reporting inter-annotator agreement rates as low as 65.7% (alpaca_eval), indicating that such tasks fundamentally lack universally accepted preference orderings. Consequently, training models with ambiguous and low-quality training data may introduce or exacerbate non-transitive preference relationships, undermining the development of stable and reliable evaluator LLMs. This phenomenon underscores the critical necessity of implementing robust filtering mechanisms to eliminate unreliable data points prior to model training.
In this paper, we present a novel graph-theoretic approach to assess and mitigate preference non-transitivity in evaluator LLMs. Our method, ELSPR, formulates multi-response pairwise comparisons as tournament graphs and systematically filters preference data that induce to overall preference non-transitivity. To quantify preference clarity, we introduce a new metric based on two-dimensional structural entropy of directed graphs. Experimental results demonstrate that models fine-tuned on cleaned data reduce preference non-transitivity by 13.78% and structural entropy by 0.0879, while achieving more robust rankings in real-world evaluation systems, as evidenced by increased standard deviations in MT-bench evaluation framework metrics. Human evaluation studies further validate our approach, confirming that data inducing to overall preference non-transitivity correlates with low inter-annotator agreement among human evaluators and poor alignment between LLM evaluations and human majority votes. These findings underscore ELSPR’s effectiveness in developing more reliable evaluation systems for LLMs. Specifically, our contributions are:
  • We introduce a graph-theoretic approach to systematically study the non-transitivity problem in preference data generated by evaluator LLM, reveal significant non-transitivity, and propose a two-dimensional structural entropy to quantify preference clarity.
  • We propose a robust filtering methodology leveraging tournament graph theory to systematically identify and eliminate preference data that induce non-transitivity in evaluator LLMs. Our experimental results demonstrate that this approach significantly reduces preference non-transitivity, decreases structural entropy, and substantially enhances the evaluation robustness of the resulting models when deployed in real-world evaluation systems.
  • We empirically demonstrate through human evaluation that data causing preference non-transitivity is inherently ambiguous and low-quality, with significantly lower inter-annotator agreement (34.4% vs. 52.6%) and lower consistency between evaluator LLM assessments and human majority votes (51.2% vs. 80.6%).
@@ 2.1 LLM-as-a-Judge and Its Non-Transitive Preferences
The prevailing LLM-as-a-Judge paradigm operates primarily through pairwise comparisons, as evidenced in established frameworks such as VicunaEval, AlpacaEval, and Arena-Hard (vicuna2023; alpaca_eval; DBLP:journals/corr/abs-2406-11939). These systems collect responses generated by various LLMs for a given set of questions and then employ advanced LLMs as judges to determine preference orders between response pairs, thus evaluating the relative performance of different models. Recent studies demonstrate that even advanced models such as GPT-4 exhibit significant preference non-transitivity when used in evaluation systems, substantially undermining the reliability of evaluation outcomes. Despite systematic exploration of six distinct prompt templates, improvements in mitigating preference non-transitivity remain limited (xu2025investigating). This issue extends to Reinforcement Learning from Human Feedback (RLHF), where non-transitive preferences significantly impair performance. Recent research suggests that this non-transitivity can hinder the learning algorithm from stably converging to the global optimum, leading to preference cycles and thus hindering the effective utilization of preference data (zhou2025extragradient).
To address this critical challenge, we systematically investigate the underlying conditions that generate non-transitivity in LLM-as-a-Judge systems and introduce a novel methodology to mitigate such inconsistencies.
Figure 2: ELSPR (Evaluator LLM training data Self-purification non-transitive Preferences via tournament graph Reconstruction) framework overview. (a) Raw preference data is collected via pairwise comparisons conducted by an advanced LLM. (b) The core analysis and filtering process: The raw data is first modeled as a tournament graph to identify cycles within SCCs. These cycles are then broken by reconstructing each SCC into a DAG based on in-degree ranking. The final global DAG serves as a rule to filter the initial raw data, separating it into a cleaned, transitively consistent training set and a discarded set of non-transitive preferences.
@@ 2.2 Data Selection for LLM Fine-tuning
Previous research (chen2024alpagasus; ge-etal-2024-clustering; li-etal-2024-one) has established that extracting high-quality subsets from synthetic training sets is crucial for effective fine-tuning. This approach not only enhances the performance of the model, but also substantially reduces the computational requirements. However, this strategy remains underexplored in the domain of LLM-as-a-Judge. Existing methodologies predominantly focus on scaling model architectures or refining engineering techniques, while neglecting the potential benefits of optimizing training data quality.
@@ 3 Methodology
This section presents a graph-theoretic analysis method using tournament graphs to analyze pairwise preferences from evaluator LLMs. The method introduces quality analysis criteria for training data, including non-transitivity detection via strongly connected components (SCCs) and preference clarity analysis via graph entropy. Additionally, a data filtering method is proposed to mitigate non-transitive preferences. Figure 2 outlines the overall framework, with implementation details provided in the following subsections.
@@ 3.1 Background
Modeling preferences generated by LLM as a tournament graph. For each question $q_{i}\in Q$, given the response set $A_{i}=\{a_{1},a_{2},...,a_{n}\}$ from $n$ LLMs, we construct a tournament graph $G_{i}=(V_{i},E_{i})$ through the following procedure: vertices: set $V_{i}=\{v_{1},v_{2},...,v_{n}\}$ corresponds to responses $A_{i}=\{a_{1},a_{2},...,a_{n}\}$. Edges: Defined by preferences generated by evaluator LLM:
@@
Here, $\mathcal{J}(a_{j},a_{k})\in\{\text{`win'},\text{`lose'}\}$ denotes the pairwise comparison result between answers $a_{j}$ and $a_{k}$, where $a_{j}$ appears before $a_{k}$ in the prompt.
Considering the common position bias in evaluator LLMs (wang-etal-2024-large-language-models-fair), we apply position swapping by comparing each response pair in both orders: $\mathcal{J}(a_{j},a_{k})$ and $\mathcal{J}(a_{k},a_{j})$. This ensures a more robust and balanced assessment (zheng2024judging_mt). If the preferences differ across orders—indicating possible position bias—a bidirectional edge is added between the corresponding vertices to represent a ‘tie’.
3.2 Quality Analysis Framework for Evaluator LLM Training Data
We introduce a framework for analyzing evaluator LLM training data quality. Leveraging directed graph representations of preference tournaments, the framework introduces two metrics: SCC analysis quantifying preference non-transitivity and two-dimensional structural entropy measuring preference clarity.
Quantifying preference non-transitivity via SCC analysis. SCCs are maximal subgraphs where any two vertices are mutually reachable through directed paths. This property provides a natural mechanism for identifying preference cycles—a direct manifestation of non-transitivity in evaluator LLM judgments. When vertex $v_{i}$ can reach vertex $v_{j}$ via a directed path, this indicates a preference chain $a_{j}\succ\cdot\succ a_{i}$. The simultaneous existence of paths from $v_{i}$ to $v_{j}$ and from $v_{j}$ to $v_{i}$ represents contradictory preference relationships, signaling a violation of transitivity.
We employ Tarjan’s algorithm (doi:10.1137/0201010) to efficiently identify SCCs in our directed graphs. Tarjan’s algorithm operates in linear time, specifically $\mathcal{O}(|V|+|E|)$, where $|V|$ and $|E|$ denote the number of vertices and edges, respectively, ensuring scalability to large graphs. Since non-transitivity requires at least three elements in a cycle, we focus on SCCs containing more than two vertices. Additionally, we exclude cases where every vertex pair shares bidirectional edges (indicating ‘tie’), as these represent consistent preference relations. For example, the pattern $(A=B)$, $(B=C)$, $(C=A)$ constitutes a valid SCC but maintains transitive preference relationships. Formally, we identify the set of non-transitive SCCs as:
$\mathcal{S}_{\text{n-t}}=\left\{S\in\text{SCCs}(G)\,\middle|\,\begin{aligned} |S|>2\land\exists v_{j},v_{k}\in S,\\ (v_{j}\leftrightarrow v_{k})\notin E(S)\end{aligned}\right\},$ (2)
where $\text{SCCs}(G)$ represents all SCC in graph $G$, $|S|$ denotes the component size, and $(v_{j}\leftrightarrow v_{k})\notin E(S)$ indicates the absence of bidirectional edges between some vertices.
To quantify the prevalence of non-transitivity across our entire training set, we compute the non-transitivity ratio:
@@ 3.2 Quality Analysis Framework for Evaluator LLM Training Data
where the numerator represents the total number of vertices in non-transitive SCCs across all questions $\mathbf{Q}$, and the denominator represents the total number of vertices across all graphs. This metric ranges from 0 to 1, with higher values indicating greater prevalence of non-transitive in the evaluator LLM’s preferences. A perfectly transitive preference system would yield $\rho_{\text{non-trans}}=0$, while completely cyclical preferences would approach $\rho_{\text{non-trans}}=1$.
@@
Analysis of preference clarity based the structural entropy of directed graph. While the SCC-based approach effectively quantifies preference non-transitivity, this metric alone is insufficient to characterize preference linearity. For instance, a dataset comprised entirely of preference ‘tie’ would exhibit perfect transitivity yet fail to establish any linear ordering. To address this limitation, we introduce directed graph entropy as a complementary measurement. Structural entropy (7456290) extends Shannon entropy (6773024) to directed graphs, providing a measure of system uncertainty and the complexity of relationships within the graph. These concepts have been widely applied across various domains  (10.1145/3616855.3635820; Duan_Chen_Liu_Liu_Yue_Li_2024; 10.1145/3660522; Hou_Zhu_Liu_Su_Xia_Wu_Xu_2025). We introduce and refine the two-dimensional structural entropy metric for directed graphs to analyze the overall clarity of preferences in evaluator LLMs.
@@
A crucial observation drives our methodology is that SCCs composed of single vertices inherently exhibit strict transitivity. Preference relations among such singleton components naturally form a linear order that does not increase preference complexity. For example, consider preferences $A\succ B$, $B\succ C$, and $A\succ C$. These form a clear linear order $A\succ B\succ C$, where each vertex constitutes its own singleton SCC. This clarity reflects the inherent transitivity of singleton SCCs. Based on this insight, we adopt SCCs as the basic community units when calculating structural entropy. To ensure accurate quantification of meaningful structural complexity, we exclude interactions between pure single-point SCCs. Instead, we retain only interactions between singleton SCCs and multi-vertex SCCs, as well as interactions between different multi-vertex SCCs. This selective approach ensures more accurate quantification of meaningful structural complexity. Consequently, preference relations closer to a linear order produce lower entropy values.
Figure 3: Cases of High and Low Structural Entropy in Preference Tournaments.
Figure 3 illustrates two contrasting scenarios that provide an intuitive explanation of the two-dimensional structural entropy in tournament graph preference relations. Figure 3 (a) depicts a high entropy case with a chaotic tournament containing multiple preference cycles resulting in complex non-transitive relationships. In contrast, Figure 3 (b) shows a low entropy case with a perfectly ordered tournament where responses form a clear linear hierarchy, producing a simple transitive structure.
Given a directed graph $G=(V,E)$ with $n=|V|$ vertices, we perform SCC decomposition resulting in $\text{SCCs}(G)=\{SCC_{1},SCC_{2},\ldots,SCC_{L}\}$, where $L$ is the total number of SCCs. For any vertex $v\in V$, we denote $d_{\text{in}}(v)$ and $d_{\text{out}}(v)$ as the in-degree and out-degree of $v$ in $G$. For any $SCC_{i}$, we define its volume as $v(SCC_{i})=\sum_{v\in SCC_{i}}d_{\text{in}}(v)$, while $v(G)$ represents the total in-degree of the entire graph.
The two-dimensional structural entropy is computed as:
$H^{2}(G)$ $=-\sum_{j=1}^{L}\frac{g_{j}}{v(G)}\log_{2}\frac{v(SCC_{j})}{v(G)}$
$-\sum_{j=1}^{L}\frac{v(SCC_{j})}{v(G)}\left(\sum_{v\in SCC_{j}}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\log_{2}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\right),$ (4)
The first term captures the entropy of the partition (inter-community complexity), while the second term captures the weighted average of entropies within each SCC (intra-community complexity). The variable $g_{j}$ represents the number of incoming edges to $SCC_{j}$ from vertices outside this component, specifically counting edges between singleton SCCs and multi-vertex SCCs, as well as edges between different multi-vertex SCCs. This value quantifies the external influence on each SCC and plays a crucial role in measuring cross-community complexity.
To facilitate meaningful comparisons across graphs of different sizes, we normalize the structural entropy. The normalized structural entropy for a graph $G$ is defined as:
$\tau(G)=\frac{H^{2}(G)}{\log_{2}n}.$ (5)
Normalization is necessary since raw entropy increases with graph size. Dividing by $\log_{2}n$ scales the entropy to a range between 0 and 1, where values closer to 0 indicate more ordered structures approaching linear hierarchy, while values closer to 1 suggest higher complexity and less clear preference ordering. The upper bound $\log_{2}n$ represents the maximum possible entropy for a graph with $n$ vertices, corresponding to a completely uniform distribution of structural information. To evaluate the overall preference clarity of an evaluator LLM across a set of questions $\mathbf{Q}$, we compute the average normalized structural entropy: